OpenStudy (anonymous):

Suppose that in one week, your company takes in $2,220 in sales at one of its cell phone stores. The price breakdown per phone is as follows: Cell phone model A4: $35 Smartphone Z20: $50 Suppose that a total of 51 phones were sold. Set up the system of equations that needs to be solved to determine how many of each type of phone were sold. Give a clear definition of the variables in the system. Solve the system of equations, showing clearly how the solution was determined, and state the results clearly in light of the real-world situation. Verify your results of the 2 linear eq by grap

OpenStudy (aripotta):

ok...cell phone model A4 will be represented as a, and smartphone Z20 will be represented as b. 35a + 50b = 2220 a + b = 51 do you understand this so far?

OpenStudy (anonymous):

Yes I do

OpenStudy (aripotta):

ok, so now let's focus on the second equation to find out what a is. to find out what a would equal, subtract b from both sides to get 51 - b. so a = 51 - b

OpenStudy (aripotta):

now we can substitute 51 - b in our first equation. 35(51 - b) + 50 b = 2220 do you know what to do now?

OpenStudy (aripotta):

50b*

OpenStudy (anonymous):

I think I do. Can you go a little further for me?

OpenStudy (aripotta):

sure. use the distributive property to distribute the 35 into the parentheses first

OpenStudy (anonymous):

Ok.

OpenStudy (aripotta):

tell me what you get so i know you're doing it lol

OpenStudy (anonymous):

So far I have 2220=35+50.

OpenStudy (aripotta):

you don't have the variables in there

OpenStudy (aripotta):

and you didn't do the distributive property

OpenStudy (anonymous):

I'm also transforming this into x and y.

OpenStudy (aripotta):

do you know how to distribute?

OpenStudy (aripotta):

the letter of the variables doesn't matter. just make sure you tell me which is which

OpenStudy (anonymous):

Okay, so I've gotten the following placed as well 2220=35*(51-y)+50*y

OpenStudy (anonymous):

Let me know how I'm doing so far.

OpenStudy (aripotta):

ok, so do know how to distribute?

OpenStudy (anonymous):

I guess I don't quite understand how to distribute?

OpenStudy (anonymous):

Here is what I've gotten so far 2220=1785+15*y; 15*y=2220-1785; 15*y=435 y=435/15 y=29

OpenStudy (aripotta):